ZO19 Introductory Physics II Main Menu Contents Grades Course Contents » HOMEwORK » Set 2 ( .) Timer □ Notes Feedback 을 Print Evaluate Two parallel plates, each of area 3.27 cm2, are separated by 4.30 mm. The space between the plates is filled with air. A voltage of 5.25 V is applied between the plates. Calculate the magnitude of the electric field between the plates. Submit Answer Tries 0/20 Calculate the amount of the electric charge stored on each plate. Submit Answer Tries 0/20 Now distilled water is placed between the plates and the capacitor is charged up again to the same voltage as before. Calculate the magnitude of charge stored on eăch plate in this case. (Use K83.0 for the dielectric constant of water.) Socon An Tries 0/20 1 Post Discussion Send Feedback
Answer #1

a)

E = V/d = 5.25 / (4.3*10^-3)

E= 1221 N/C

========

b)

E = q / (A * epsilon_not)

q = 1221* 3.27*10^-4* 8.85*10^-12

q = 3.53*10^-12 C

========

c)

Q = k q = 83 × 3.53*10^-12

Q = 2.93*10^-10 C

========

Comment in case any doubt, will reply for sure. Goodluck